OpenStudy (anonymous):

Can you simplify this further? *Using the quadratic formula to factor x^2-6x-4* The answer is supposed to be 6(+-)√53 / 2 ,but can't you simplify by 2? 6 and 2 go into 2. Then could it be 3(+-)√68/1 or no?

hero (hero):

Hi neverforgettovisitme

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

6(+-)√53 / 2 is simplified

hero (hero):

:D

OpenStudy (anonymous):

answer is is \[3\pm\sqrt{13}\]

OpenStudy (anonymous):

wait how is it simplified? in another problem there was a problem where you divided by 3 because it was a common factor

OpenStudy (anonymous):

wait i'll go look for it

OpenStudy (anonymous):

-9x^2-9x+6

OpenStudy (anonymous):

no you cannot because the number inside the radical contains no perfect squares. this is the best you can do

OpenStudy (anonymous):

later in the problem the guy got 9(+-)3√33 --------- -18 then he simplified by 3

OpenStudy (anonymous):

khanacademy

OpenStudy (anonymous):

for \[-9x^2-9x+6=0\] you can start with \[3x^2+3x-2=0\] and reduce before you start

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

hold the phone solution to \[ x^2-6x-4=0\]is not what you wrote. it is what tomas A wrote

OpenStudy (anonymous):

\[\frac{9\pm3\sqrt{33}}{-18}=\] you can simplify this since \[\frac{3(3\pm\sqrt{33})}{-6\cdot3}=\] \[\frac{1(3\pm\sqrt{33})}{-6\cdot1}=\]

OpenStudy (anonymous):

\[x^2-6x-4=0\] \[x^2-6x=4\] \[(x-3)^2=4+9=13\] \[x-3=\pm\sqrt{13}\] \[x=3\pm\sqrt{13}\]